On January 4,2021, Tamara Bakery paid ₱30 million for 1 million shares of jade Company ordinary shares. The investment represents a 20% interest in the net assets of jade and gave Tamara the ability to exercise significant influence over Jade’s operations. Tamara received dividends of ₱1.00 per share on December 15,2021 and jade reported net income of ₱8 million for the year ended December 31,2021. The market value of Jade’s ordinary shares at December 31,2021 was ₱32 per share, on the purchase date, the book value of Jade’s net assets was ₱120 million and the fair market value of Jade’s depreciable assets with an average remaining useful life of six years exceeded their book value by ₱6 million. The remainder of the excess of the cost of the investment over the book value of net assets purchased was attributable to goodwill.


1. How much is the implied goodwill from acquisition?


1
Expert's answer
2021-06-14T14:02:36-0400

Solution:

Implied goodwill: Carry out the two-step process:

The first step is to compare the fair value of the assets with the book value of the assets. If the fair value is lower, the goodwill is impaired.

Derive goodwill from the acquisition:

Goodwill = purchase price less net book value of assets less fair value adjustments

Purchase price = 30M

Net book value of assets ="120M\times 20\% = 24M"

Excess purchase price = 30M – 24M = 6M

Fair value adjustments = Fair value of assets – book value of assets

Fair value of assets = "120M + 6M = 126M\times 20\% = 25.2M"

Net Book value of assets = 24M

Fair value adjustments = 25.2M – 24M = 1.2M

Goodwill = 30M – 24M – 1.2M = 4.8M

Goodwill = 4.8M


Book value of assets inclusive of goodwill = 24M + 4.8M = 28.8M

Fair value of assets = 25.2M

Thus, the fair value of the reporting unit is less than its book value, therefore, an impairment has occurred.

 

The second step involves comparing the fair value of the assets with the book value net of goodwill and the difference is the implied goodwill amount.

The fair value of assets = 25.2M

Book value of assets net of goodwill = 28.8M – 4.8M = 24M

Implied goodwill = 25.2M – 24M = 1.2M

Implied goodwill from acquisition = 1.2M